(a) Prove that for all integer n à ¢ ‰ ¥ 3, P (n +1,3) - P(n,3) = 3P(n,2) b)Prove that n.P(n-1,n-1) = P(n,n) c)show that C(n+1,k) = C (n,k-1) + C(n,k)
1
Expert's answer
2021-10-21T14:04:08-0400

(a)


"P(n+1,3)=\dfrac{(n+1)!}{(n+1-3)!}=(n+1)(n)(n-1)"

"P(n,3)=\dfrac{(n)!}{(n-3)!}=n(n-1)(n-2)"

"P(n,2)=\dfrac{(n)!}{(n-2)!}=n(n-1)"

Then for "n\geq3"


"P(n+1,3)-P(n,3)"

"=(n+1)(n)(n-1)-n(n-1)(n-2)"

"=n(n-1)(n+1-n+2)"

"=3n(n-1)=3P(n,2)"

(b)


"P(n-1,n-1)=\dfrac{(n-1)!}{(n-1-n+1)!}=(n-1)!"

"P(n,n)=\dfrac{(n)!}{(n-n)!}=n!"

Then


"nP(n-1, n-1)=n(n-1)!=n!=P(n, n)"

(c)


"C(n+1,k)=\dbinom{n+1}{k}=\dfrac{(n+1)!}{k!(n+1-k)!}"

"C(n,k-1)=\dbinom{n}{k-1}=\dfrac{n!}{(k-1)!(n-k+1)!}"

"C(n,k)=\dbinom{n}{k}=\dfrac{n!}{k!(n-k)!}"

Then


"C(n,k-1)+C(n,k)"

"=\dfrac{n!}{(k-1)!(n-k+1)!}+\dfrac{n!}{k!(n-k)!}"

"=\dfrac{n!(k+n-k+1)}{k!(n-k+1)!}=\dfrac{n!(n+1)}{k!(n-k+1)!}"

"=\dfrac{(n+1)!}{k!(n+1-k)!}=C(n+1,k)"